# MLRC-Bench: Can Language Agents Solve Machine Learning Research Challenges?

### v2

@@ 1 Introduction
Evaluating large language model (LLM) research agents [2, 32, 36] has so far been restricted to two directions. One involves tasking the agent with end-to-end scientific discovery—proposing a research idea, writing implementation code, running experiments, and eventually producing a full paper as done by AI Scientist [36]. One issue with such evaluation is the lack of reliable baseline method that enables objective evaluation of the proposed approach. The second direction, on the other hand, evaluates the agent’s ability to produce code that solves a Kaggle-style machine learning (ML) engineering competition, skipping idea proposal and paper writing altogether [20, 7]. While evaluation in this case is straightforward, this setup rarely demands genuine research novelty beyond existing methods. Consequently, neither of these setups paints a full picture of whether LLM research agents can design research ideas that are both novel and effective, which we aim to address here.
@@
We curate MLRC-Bench starting with 7 competition tasks as shown in Table 1. We pick tasks that involve novel and high-impact problems, spanning areas including LLM safety, multimodal perception, and few-shot learning. Our experimental findings reveal that even the best-performing tested LLM agents, such as gemini-exp-1206 [42] under the MLAB [20] scaffolding, closes only 9.3% of the gap between baseline and top human participant score. Additionally, our analysis highlights a poor correlation between the novelty judged by LLM and practical effectiveness of agents’ solutions, questioning the reliability of LLM-as-a-judge for research idea evaluation. These results underscore the limitations of current AI research agents in generating and implementing innovative ML solutions, providing a crucial benchmark for future advancements.
Our contributions can be summarized as below: 1) We introduce MLRC-Bench, a dynamic benchmark suite curated from ML conference competitions, featuring open research problems that are both impactful and objectively measurable, and that demand the development of novel methodologies; 2) We conduct large-scale, objective evaluations for a wide array of frontier LLMs with representative agent scaffoldings, highlighting their inability to propose and implement innovative solutions with notable performance gains; 3) We pinpoint the flaws in subjective evaluations of LLM-based research agents, by showing that the LLM-judged idea novelty is misaligned with empirical effectiveness.
@@ 2 Related Work
While there are recent benchmarks that focus on code generation in machine learning domain, they do not always require methodological innovation. Works like MLAgentBench [20] and MLE-Bench [7] evaluate agents on Kaggle-style ML tasks but prioritize code implementation over novel research contributions. MLE-Bench requires the final submission to be a CSV file, limiting the data modality. Broader benchmarks such as ScienceAgentBench [13] and DiscoveryBench [38] span multiple scientific domains but lack granularity for ML-specific challenges, while CHIME [16] and OpenD5 [14] target auxiliary tasks like literature review or hypothesis generation. DSBench [27] and AAAR-1.0 [34] extend evaluations to data science and general R&D workflows but still fall short of addressing cutting-edge ML research innovation. RE-Bench [50] and MLGym [39] provide collections of ML research task environments, but their problems either fail to represent most recent research directions (e.g., image classification with CIFAR-10 [30]), or only cover a narrow range of research domains. For example, 6 out of 7 tasks in RE-bench are related to language modeling. Besides, existing benchmarks often fail to specify computation constraints (e.g. runtime and GPU memory limit), which are important to encourage efficient yet effective solutions. The differences between our benchmark and existing work on automating ML research workflow with LLM agents are summarized in Table 3 of Appendix B.
@@
Unlike DiscoPOP [35] and DA-Code [21], which focus on function-level coding or data science, MLRC-Bench requires repository-level code comprehension and generation, thus better reflecting the capabilities required to work with realistic and complex research codebases. This design enables substantial project-level coding tasks, such as generating and editing multiple files [37] and reusing existing utility functions [33]. Moreover, repository-level design supports multi-agent collaboration, where specialized agents for tasks like literature survey, idea brainstorming, code writing, and performance evaluation can iteratively work together [19].
@@
Automated end-to-end research workflows like The AI Scientist [36, 52] and MLR-Copilot [32] rely largely on subjective reviews of papers or research proposals for evaluating success. In parallel, ResearchAgent [2] iteratively refines ideas through multi-agent feedback, and Chain-of-Idea-Agent [31] organizes literature into progressive chains to stimulate ideation. However, it remains unclear how subjectively evaluated “novel” ideas translate into actual performance gains. In contrast, we explicitly investigate how such subjective assessments of novelty or idea quality align—or fail to align—with measurable performance improvements.
@@ 3 MLRC-Bench
3.1 Task Selection
MLRC-Bench prizes high-quality competitions that are both non-trivial and reproducible. To form our dataset, we screen competitions held at recent machine learning, natural language processing, data mining, and computer vision conferences or workshops using the following criteria:
  • Novel Research-Focused: The tasks should require genuine methodological innovation, rather than being solvable through purely brute-force or superficial engineering approaches, such as exhaustive search for hyperparameters or features without any theoretical motivation or problem understanding.
  • Non-Trivial: The problem must involve complexity so that it will not be solved by simply applying standard ML algorithms, e.g., calling the XGBoost classifier [11] on a new dataset or prompt engineering with LLMs.
  • Feasible: Starter code, data splits, and evaluation procedures must be publicly available so that researchers, either human or agentic AI, can reproduce the experiments while keeping computational costs manageable.
The current version of MLRC-Bench comprises seven tasks adapted from existing competitions. [^1] While numerous competitions have been hosted at recent ML/AI conferences, only a limited subset was selected due to factors such as certain challenges being considered solved with the rapid advancement of foundation models, missing or irreproducible evaluation data/code, qualitative rather than quantitative evaluation, insufficient recency, or limited emphasis on algorithmic innovation. These tasks represent a diverse landscape of applied ML research. As shown in Table 1, topics range from LLM safety to multimodal perception, ensuring that benchmarks stress multiple facets of algorithmic creativity. Besides, these tasks are cutting-edge. We select competitions released in recent years, whose problems remain actively researched, preventing trivial or purely off-the-shelf solutions. In addition, we have the following two considerations for MLRC-Bench design.
Continual Updates.
MLRC-Bench is designed to evolve over time. We will continue to expand the benchmark with tasks from recent conference competitions and retire older ones once model performance saturates. This ensures the benchmark remains aligned with the frontier of machine learning research, enabling continuous tracking of agent progress on fresh and unsolved challenges. In parallel, we have released templates and detailed instructions to facilitate community contributions. [^2] https://tinyurl.com/MLRC-Task-Template All competitions follow a uniform and standardized code structure (detailed in Section 3.2) to streamline research and accelerate progress on LLM agents.
Data Contamination Mitigation.
Although most academic competitions require top teams to submit solution reports, the actual implementation code may not be publicly available, reducing the risk of such code appearing in the pretraining data of LLMs. As previously discussed, we will regularly update the benchmark with suitable recent competitions, ensuring that it assesses an agent’s genuine research capabilities rather than its capacity for mere data retrieval and replication.
Computational Constraints.
MLRC-Bench specifies explicit computational constraints, such as runtime and GPU memory limits, closely mirroring real-world competition scenarios. These standardized constraints ensure fairness by preventing agents from leveraging excessive computational resources to gain undue advantage, thereby incentivizing efficient and innovative solutions.
Figure 1: Overview of MLRC-Bench and the evaluation pipeline.
@@ 3.2 Task Environment
  • Task Description. A detailed account of the research problem, including essential terminology, data format, desired model outputs, and constraints (e.g., limitations of model size or training time).
@@ 3.4 Evaluation Protocol
Table 2: For each competition and agent, we report the test-phase relative improvement to the human solution. Best performing agent in each task is highlighted in bold. Our results indicate that providing additional ideas, whether sourced from AI or humans, does not consistently yield performance improvements. The best-performing configuration—gemini-exp-1206 under MLAB—achieves only 9.3% of the human-level improvement over baseline on average, underscoring the inherent difficulty of these research tasks. See Table 4 in Appendix D for absolute improvements to baseline solution.
@@
4 Experiments and Results
4.1 Agent Scaffolding Comparison
In addition to allow agents to directly propose and implement ideas, we investigate whether providing AI-generated or human-sourced ideas can enhance agent performance. Due to computational cost limits, we follow the practice of MLE-Bench [7] to evaluate a commonly used model for agentic tasks, GPT-4o [22], under three scaffolding configurations:
  • MLAB: We chose a general-purpose MLAB [20] as our primary agent for evaluation.555 Another candidate agent framework, AIDE [24] is tailored for Kaggle-style tasks with single-file solutions and tabular outputs, making it less generalizable to our repository-level coding with diverse output formats (e.g., images, tensors). Its tree-search-based design with performance-based rewards is promising but requires careful adaptation, which we leave to future work. It is a ReAct-style [53] agent that alternates between thinking (such as reflection, research planning, fact-checking) and taking actions (file system operations or Python script execution) to implement methods.
  • CoI-Agent Idea + MLAB: We augment MLAB with ideas generated by Chain-of-Ideas (CoI) [31], an LLM-based agent that structures relevant literature into progressive chains to enhance ideation. We choose a strong model, OpenAI’s o1 [23], as the backbone for the ideation agent to generate more creative ideas.
  • Human Idea + MLAB: To investigate whether agents can achieve notable performance gain given the right direction to work on, we provide MLAB with reliable human ideas extracted from state-of-the-art papers and top-performing competition participants’ reports.
For all tasks, MLAB agents are allowed a maximum of 50 steps and 5 hours per trial, except for Rainfall Prediction, where the limits are extended to 100 steps and 10 hours due to the longer baseline training time. As shown in Table 2, incorporating additional ideas—whether generated by AI or proposed by humans—does not consistently improve performance. This underscores the challenges agents face not only in generating high-quality ideas but also in effectively implementing them, even when the ideas originate from humans.
@@ 4.2 Model Comparison
Taking MLAB as our major scaffold, we evaluate five prominent LLMs: Claude 3.5 Sonnet v2 [1], gemini-exp-1206 [42] [^6] Gemini-exp-1206 is renamed to Gemini 2.0 Pro afterwards., Llama 3.1 405B Instruct [15], o3-mini-high [41] and GPT-4o (2024-11-20) [22]. The results in Table 2 show varying success rates across models and tasks. Among all models, Gemini-exp-1206 performs the best overall, closing 9.3% of the gap between baseline and top human participant scores. Claude 3.5 Sonnet V2 performs the best on most tasks but fails significantly on machine unlearning. A case study of its final solution (Appendix C.1) suggests that the failure stems from treating the removal of unwanted data and the preservation of useful knowledge as separate objectives, rather than jointly optimizing both to maintain model performance. Furthermore, our cost-effectiveness evaluation in Appendix C.2 identifies Llama 3.1 405B with MLAB as a Pareto-optimal setup balancing API cost and performance.
Table 2 also reveals that agent solutions’ performance gains remain modest compared to human solutions in many cases, if not degrading baseline performance. There are, however, a few notable exceptions. For instance, MLAB (gpt-4o) achieves a score of 47.5 on the rainfall prediction task, likely because similar solutions (e.g., variants of U-Net [43]) are readily available online. In the backdoor-trigger task, the baseline GCG method [56] performs poorly—essentially making random predictions—thereby lowering the bar for agents to surpass it with more meaningful solutions. This substantial gap highlights the current limitations of AI agents in generating novel, effective methods, underscoring the need for further advances to match or surpass human-led research efforts.
Figure 2: We track the percentages of changes of performance, runtime, and lines of code compared to baseline across iterative refinement of implementations within a trial of LLM-based MLAB agent on the development set. Performance improvement is the higher the better, while increased runtime and lines of code are the lower the better. These figures show the averaged metrics across all tasks. For results breakdown on each task, please refer to Figure 9 and 10 in Appendix D. The takeaway is that agents tend to over-refine their solutions over time, leading to increased complexity and runtime without proportional performance gains.
4.3 Solution Development Analysis
Figure 2 illustrates how performance, runtime, and code complexity evolve as agents iteratively refine their implementations within a single trial. Three key trends emerge: (1) GPT-4o and Claude gradually improve their performance through refinement, while other models plateau after a few iterations; (2) runtime consistently increases, probably because models are exploring more complicated solutions over time, which may naturally conflate with better solutions; and (3) code size expands over time, reflecting increasingly complex solutions that do not yield proportional performance gains. Together, these trends suggest that agents tend to over-refine their solutions, resulting in more complex and time-consuming implementations without further performance improvements.
We further analyze the agent traces (gemini-exp-1206 with MLAB) in Appendix F to understand its behavioral patterns and limitations. Two key insights emerge. First, a significant portion of action errors—11.5% of all steps—stem from incorrect tool arguments, where the model either hallucinates or misidentifies expected parameter names. This highlights the need to improve the agent’s tool-usage capabilities. Second, the agent was able to fix only 17.2% of the errors encountered during code execution, revealing the challenge of self-debugging in complex, repository-level ML codebases [12].
@@ 4.3 Solution Development Analysis
4.4 Subjective Evaluation with LLM-as-a-Judge
Our benchmark also facilitates the investigation on whether LLM-as-a-judge style evaluations can reliably assess the quality of research ideas by comparing subjective (here as judged by o1) and objective metrics. As shown in Figure 1, we first prompt an LLM to explain each implementation’s underlying idea. [^7] We prompt the MLAB agent to include detailed comments in the code to enable faithful explanations. We choose OpenAI’s o1 [23] model here because of its superior reasoning and coding capabilities. We then apply the rubric from prior work [2] to have the o1 model assign a 1–5 Likert score on five dimensions: validity, clarity, rigorousness, generalizability, and innovativeness. These scores are all the higher the better. The prompts used for this evaluation are shown in Appendix E. Furthermore, we examine how the presence of code influences the assessments through two settings. (1) Without Code, in which the LLM judges only access the task description and the proposed idea; and (2) With Code, in which the judges also see the code implementation. We then compute Spearman’s correlation [46] for each pair of objective and subjective metrics, using data from all valid implementations that include test-phase scores.

### v3

@@ 1 Introduction
Evaluating large language model (LLM) research agents [2, 35, 40] has so far been restricted to two directions. One involves tasking the agent with end-to-end scientific discovery: proposing a research idea, writing implementation code, running experiments, and eventually producing a full paper as done by AI Scientist [40]. One issue with such evaluation is the lack of reliable baseline method that enables objective evaluation of the proposed approach. The second direction, on the other hand, evaluates the agent’s ability to produce code that solves a Kaggle-style machine learning (ML) engineering competition, skipping idea proposal and paper writing altogether [22, 7]. While evaluation in this case is straightforward, this setup rarely demands genuine research novelty beyond existing methods. Consequently, neither of these setups paints a full picture of whether LLM research agents can design research ideas that are both novel and effective, which we aim to address here.
@@
We curate MLRC-Bench starting with 7 competition tasks in Table 1. We pick tasks that involve novel and high-impact problems, spanning areas including LLM safety, multimodal perception, and few-shot learning. Our experimental findings reveal that even the best-performing tested LLM agents, such as gemini-exp-1206 [46] under the MLAB [22] scaffolding, closes only 9.3% of the gap between baseline and top human participant score. Additionally, our analysis highlights a poor correlation between the novelty judged by LLM and practical effectiveness of agents’ solutions, questioning the reliability of LLM-as-a-judge for research idea evaluation. These results underscore the limitations of current AI research agents in generating and implementing innovative ML solutions, providing a crucial benchmark for future advancements.
Our contributions can be summarized as below:
  • We introduce MLRC-Bench, a dynamic benchmark suite curated from ML conference competitions, featuring open research problems that are both impactful and objectively measurable, and that demand the development of novel methodologies;
  • We conduct large-scale, objective evaluations for a wide array of frontier LLMs with representative agent scaffoldings, highlighting their inability to propose and implement innovative solutions with notable performance gains;
  • We pinpoint the flaws in subjective evaluations of LLM-based research agents, by showing that the LLM-judged idea novelty is misaligned with empirical effectiveness.
@@ 2 Related Work
AI Scientist [40] LLM & Human Judge Paper
Can LLMs Generate Novel Research Ideas? [49] Human Judge Idea Proposal
DiscoPOP [39] Performance -Based Function-Level Code
MLAgentBench [22] $\sim$ Performance -Based Single-Script Code
MLE-Bench [7] $\sim$ Performance -Based Single-Script Code
MLGym-Bench [43] $\sim$ Performance -Based Single-Script Code
RE-Bench [59] Performance -Based Single-Script Code
MLRC-Bench (Ours) Performance -Based Repository -Level Code
@@
Scientific discovery in machine learning typically includes four main stages: Problem Identification, where gaps in existing methods are recognized; Method Proposal, which introduces a new approach to address the issue; Experiment Design, involving the selection of datasets, baselines, and metrics for evaluation; and Code Implementation, where the method is realized through executable code. While prior work [40, 49] covers Problem Identification and Experiment Design, evaluation could be subjective based on idea proposal or final paper. Instead, MLRC-Bench focuses on the critical stages of proposing and implementing novel methods, enabling objective performance assessment. While there are recent benchmarks that focus on code generation in machine learning domain, they do not always require methodological innovation. Works like MLAgentBench [22] and MLE-Bench [7] evaluate agents on Kaggle-style ML tasks but prioritize code implementation over novel research contributions. MLE-Bench requires the final submission to be a CSV file, limiting the data modality. Broader benchmarks such as ScienceAgentBench [13] and DiscoveryBench [42] span multiple scientific domains but lack granularity for ML-specific challenges, while CHIME [18] and OpenD5 [15] target auxiliary tasks like literature review or hypothesis generation. DSBench [29] and AAAR-1.0 [38] extend evaluations to data science and general R&D workflows but still fall short of addressing cutting-edge ML research innovation. RE-Bench [59] and MLGym [43] provide ML research task environments but mostly cover outdated or narrow domains (e.g., CIFAR-10 [32]), with six of seven RE-Bench tasks focusing on language modeling. As its tasks are manually curated by experts, RE-Bench is difficult to update and often lags behind emerging research trends. In contrast, MLRC-Bench sources tasks directly from ML conference competitions, ensuring continual inclusion of cutting-edge problems. Moreover, while RE-Bench evaluates single-script solutions, our benchmark features repository-level coding to better reflect real-world research workflows. Besides, existing benchmarks often fail to specify computation constraints (e.g. runtime and GPU memory limit), which are important to encourage efficient yet effective solutions.
@@
Unlike DiscoPOP [39] and DA-Code [23], which focus on function-level programming or data-science workflows, MLRC-Bench targets repository-level code comprehension and generation, more faithfully reflecting the skills needed to navigate complex research codebases. This setup supports advanced development behaviors, such as generating and editing multiple interdependent files [41] and reusing existing utilities [36]. Consequently, many ML-agent frameworks [26, 53, 37] built for single-file solutions struggle with the richer challenges posed by our benchmark. While general-purpose coding agents like MLAB [22], OpenHands [58], and RepoMaster [56] can handle repository-level tasks and produce valid implementations, they still find it difficult to iteratively optimize algorithmic effectiveness. Moreover, the repository-level design enables multi-agent collaboration, where specialized agents for literature review, idea generation, coding, and evaluation can jointly refine solutions [21, 17, 65, 48, 62].
@@
Automated end-to-end research workflows like The AI Scientist [40, 61] and MLR-Copilot [35] rely largely on subjective reviews of papers or research proposals for evaluating success. In parallel, ResearchAgent [2] iteratively refines ideas through multi-agent feedback, and Chain-of-Idea-Agent [34] organizes literature into progressive chains to stimulate ideation. However, it remains unclear how subjectively evaluated “novel” ideas translate into actual performance gains. In contrast, we explicitly investigate how such subjective assessments of novelty or idea quality align or fail to align with measurable performance improvements. The differences between our benchmark and existing work on automating ML research workflow with LLM agents are presented in Table 2.
Figure 1: Overview of MLRC-Bench and its evaluation pipeline. MLRC-Bench standardizes ML conference competitions into an agent-agnostic framework featuring repository-level code execution under compute constraints. Its evaluation relies on objective metrics (effectiveness, efficiency, simplicity) while using subjective LLM-judge scores only to analyze their correlation with objective metrics for assessing LLM-judge reliability (Section 4.3).
@@ 3.1 Task Environment
  • Task Description. A detailed description of the research problem, including essential terminology, data format, desired model outputs, and constraints (e.g., limitations of model size or training time).
@@
3.2 Task Selection
MLRC-Bench prizes high-quality competitions that are both non-trivial and reproducible. To form our dataset, we screen competitions held at recent machine learning, natural language processing, data mining, and computer vision conferences or workshops using the following criteria: 1) Novel Research-Focused: The tasks should require genuine methodological innovation, rather than being solvable through purely brute-force or superficial engineering approaches, such as exhaustive search for hyperparameters or features without any theoretical motivation or problem understanding. 2) Non-Trivial: The problem must involve complexity so that it will not be solved by simply applying standard ML algorithms, e.g., calling the XGBoost classifier [11] on a new dataset or prompt engineering with LLMs. 3) Feasible: Starter code, data splits, and evaluation procedures must be publicly available so that researchers, either human or agentic AI, can reproduce the experiments while keeping computational costs manageable.
The current version of MLRC-Bench comprises seven tasks adapted from existing ML research competitions. [^1] While numerous competitions have been hosted at recent ML/AI conferences, only a limited subset was selected due to factors such as certain challenges being considered solved with the rapid advancement of foundation models, missing or irreproducible evaluation data/code, qualitative rather than quantitative evaluation, insufficient recency, or limited emphasis on algorithmic innovation. These tasks span a diverse landscape of applied ML research, covering topics from LLM safety to multimodal perception (Table 1), and are carefully curated to reflect unsolved, high-impact challenges that demand genuine algorithmic creativity rather than incremental tuning or ensembling. Rather than emphasizing breadth across numerous tasks [7], MLRC-Bench prioritizes depth and research-grade complexity, where solving even a single task signifies meaningful scientific progress.
In addition, we have the following three considerations for MLRC-Bench design. Continual Updates. MLRC-Bench evolves with the field by adding new ML conference competitions and retiring tasks where performance has saturated, ensuring alignment with frontier research. We also provide standardized templates and contribution guidelines to encourage community expansion. [^2] https://tinyurl.com/MLRC-Task-Template Data Contamination Mitigation. Moreover, since competition solutions are typically shared only as summary reports rather than full implementation code, their content rarely appears in LLM pretraining data, minimizing contamination. Regular updates further ensure that the benchmark evaluates genuine research ability rather than memorized solutions. Computational Constraints. Finally, each task enforces explicit runtime and GPU memory limits that mirror real-world competition settings, ensuring fairness and encouraging efficient, resource-conscious methods.
@@ 3.4 Evaluation Protocol
4 Experiments and Results
4.1 Agent Scaffolding Comparison
In addition to allowing agents to directly propose and implement ideas, we investigate whether providing AI-generated or human-sourced ideas can enhance agent performance. Due to computational cost constraints, we follow the practice of MLE-Bench [7] to evaluate a commonly used model for agentic tasks, GPT-4o [24], under three scaffolding configurations:
  • MLAB: We adopt the general-purpose MLAB framework [22] as our primary agent for evaluation. [^5] While existing ML agent frameworks, including AIDE [26], SELA [14], and AutoML-Agent [55], demonstrate strong performance on Kaggle-style tasks that yield single-file solutions, their design assumptions differ substantially from our repository-level coding setup, necessitating careful adaptation for effective use. MLAB is a ReAct-style [63] agent that alternates between reasoning steps (e.g., reflection, research planning, fact-checking) and actions (e.g., file operations or Python execution) to implement machine learning methods. Notably, MLAB does not include any built-in web search capability; it operates entirely over local resources such as code files, the Python runtime, and its internal memory.
  • CoI-Agent Idea + MLAB: We augment MLAB with ideas generated by Chain-of-Ideas (CoI) [34], an LLM-based ideation agent that structures relevant literature into progressive reasoning chains. CoI-Agent is equipped with web-based tools, including access to the Semantic Scholar API for retrieving up-to-date research papers. We use OpenAI’s o1 [25] model as its backbone to encourage more creative and literature-grounded ideation. By evaluating CoI-Agent-generated ideas with MLAB, we effectively study how agents can leverage web-based retrieval to tackle machine learning research challenges.
  • Human Idea + MLAB: To test whether agents can achieve stronger performance when given high-quality conceptual guidance, we provide MLAB with human-curated ideas manually extracted from state-of-the-art papers or top-performing participants’ reports.
For all tasks, MLAB agents are limited to 50 steps and 5 hours per trial, except for Rainfall Prediction, which allows 100 steps and 10 hours to match the longer training time of the official baseline. This ensures fair comparison, as the baseline requires more epochs to converge due to a larger dataset. As shown in Table 3, incorporating additional ideas, whether generated by AI or proposed by humans, does not consistently improve performance. This underscores the challenges agents face not only in generating high-quality ideas but also in effectively implementing them, even when the ideas originate from humans.
Table 3: For each competition and agent, we report the test-phase relative improvement to the human solution. Best performing agent in each task is highlighted in bold. Our results indicate that providing additional ideas, whether sourced from AI or humans, does not consistently yield performance improvements. The best-performing configuration, gemini-exp-1206 under MLAB, achieves only 9.3% of the human-level improvement over baseline on average, underscoring the inherent difficulty of these research tasks. See Table 4 in Appendix B for absolute improvements to baseline solution.
@@ 4.2 Model Comparison
Taking MLAB as our major scaffold, we evaluate five prominent LLMs: Claude 3.5 Sonnet v2 [1], gemini-exp-1206 [46] [^6] Gemini-exp-1206 is renamed to Gemini 2.0 Pro afterwards., Llama 3.1 405B Instruct [16], o3-mini-high [45] and GPT-4o (2024-11-20) [24]. The results in Table 3 show varying success rates across models and tasks. Among all models, Gemini-exp-1206 performs the best overall, closing 9.3% of the gap between baseline and top human participant scores. Claude 3.5 Sonnet V2 performs the best on most tasks but fails significantly on machine unlearning. A case study of its final solution (Appendix C.1) suggests that the failure stems from treating the removal of unwanted data and the preservation of useful knowledge as separate objectives, rather than jointly optimizing both to maintain model performance.
Table 3 also reveals that agent solutions’ performance gains remain modest compared to human solutions in many cases, if not degrading baseline performance. There are, however, a few notable exceptions. For instance, MLAB (gpt-4o) achieves a score of 47.5 on the rainfall prediction task, likely because similar solutions (e.g., variants of U-Net [47]) are readily available online. In the backdoor-trigger task, the baseline GCG method [68] performs poorly, where it essentially makes random predictions, thereby lowering the bar for agents to surpass it with more meaningful solutions. This substantial gap highlights the current limitations of AI agents in generating novel, effective methods, underscoring the need for further advances to match or surpass human-led research efforts.
@@
4.3 Reliability of Subjective Evaluations with LLM-as-a-Judge
Our benchmark enables an investigation of whether LLM-as-a-judge evaluations [33, 66] can reliably assess the quality of research ideas by comparing subjective judgments against objective performance. As illustrated in Figure 1, we first prompt an LLM to explain each implementation’s underlying idea. [^7] We prompt the MLAB agent to include detailed comments in the code to enable faithful explanations. Following a peer-reviewed rubric [2], an LLM then assigns 1–5 Likert scores for the explained ideas on five dimensions: validity, clarity, rigorousness, generalizability, and innovativeness. These scores are all the higher the better. The prompts used for this evaluation are shown in Appendix E. We adopt OpenAI’s o1 [25] model as both the idea explainer and the judge for its strong reasoning and coding capabilities. The subjective scores serve only to analyze potential biases in LLM-based evaluation; they are not part of the benchmark evaluation protocol.
Furthermore, we examine how the presence of code influences the assessments through two settings. (1) Without Code, in which the LLM judges only access the task description and the proposed idea; and (2) With Code, in which the judges also see the code implementation. We then compute Spearman’s correlation [51] for each pair of objective and subjective metrics, using data from all valid implementations that include test-phase scores.